# Semi-Supervised Deep Sobolev Regression: Estimation, Variable Selection and Beyond

### v1

@@ Appendix B Proofs of Results in Section 3
$|{\mathscr{B}}(f,g)|$ $\leq(\lambda\vee\zeta^{-1/2})\|f\|_{H^{1}(\nu_{X})}\|g\|_{H^{1}(\nu_{X})},\quad f,g\in H^{1}(\nu_{X}),$
${\mathscr{B}}(f,f)$ $\geq(\lambda\wedge\kappa^{-1/2})\|f\|_{H^{1}(\nu_{X})},\quad f\in H^{1}(\nu_{X}).$
Further, since that $f_{0}\in L^{2}(\mu_{X})$, the functional $F:H\rightarrow{\mathbb{R}},~{}g\mapsto(f_{0},g)_{L^{2}(\mu_{X})}$ is bounded and linear. Then according to Lax-Milgram theorem (Evans, 2010, Theorem 1 in Chapter 6.2), there exists a unique solution $f^{\lambda}\in H^{1}(\nu_{X})$ to the varitional problem (23). This completes the proof of the uniqueness. See Grisvard (2011, Theorem 2.4.2.7) for the proof of the higher regularity of the solution. ∎
Appendix C Proofs of Results in Section 4
Proof of Lemma 4.1.
The proof uses some important results that will be verified in the proof of Lemma 3.1 and thus whose proofs are omitted here. Recall the population excess risk $R(f)$ and the empirical excess risk $\widehat{R}_{{\mathcal{D}}}(f)$ defined in the proof of Lemma 3.1. We further define the regularized excess risk and regularized empirical risk as
$R^{\lambda}(f)=R(f)+\lambda\|\nabla f\|_{L^{2}(\nu_{X})}^{2}\quad\text{and}\quad\widehat{R}_{{\mathcal{D}}}^{\lambda}(f):=\widehat{R}_{{\mathcal{D}}}(f)+\lambda\|\nabla f\|_{L^{2}(\nu_{X})}^{2}.$
@@ Appendix C Proofs of Results in Section 4
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R^{\lambda}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}\lesssim\inf_{f\in{\mathcal{F}}}R^{\lambda}(f)+(B_{0}^{2}+\sigma^{2})\frac{\operatorname{VCdim}({\mathcal{F}})}{n\log^{-1}(n)}.$ (24)
Firstly, it follows from (37) and Lemma A.4 that
$R^{\lambda}(\widehat{f}_{{\mathcal{D}}}^{\lambda})-2\widehat{R}_{{\mathcal{D}}}^{\lambda}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\leq R(\widehat{f}_{{\mathcal{D}}}^{\lambda})-2\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\leq cB_{0}^{2}\frac{\operatorname{VCdim}({\mathcal{F}})}{n\log^{-1}(n)},$ (25)
where $c$ is an absolute positive constant. It remains to consider the regularized empirical risk. According to (1), we have that for each fixed $f\in L^{\infty}(\Omega)$
$\widehat{L}_{{\mathcal{D}}}^{\lambda}(\widehat{f}_{{\mathcal{D}}}^{\lambda})=\widehat{R}_{{\mathcal{D}}}^{\lambda}(\widehat{f}_{{\mathcal{D}}}^{\lambda})-\frac{2}{n}\sum_{i=1}^{n}\xi_{i}(\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})-f_{0}(X_{i}))+{\mathbb{E}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}\Big{]}.$
Using (41) and Lemma A.4 and taking expectation with respect to ${\mathcal{D}}\sim\mu^{n}$ yields that for each $f\in{\mathcal{F}}$
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{R}_{{\mathcal{D}}}^{\lambda}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}$ $\leq{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{L}_{{\mathcal{D}}}^{\lambda}(f)\Big{]}+2{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})\Big{]}-{\mathbb{E}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}\Big{]}$
$=R^{\lambda}(f)+\frac{1}{2}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+c(B_{0}^{2}+\sigma^{2})\frac{\operatorname{VCdim}({\mathcal{F}})}{n\log^{-1}(n)},$
@@
$\widehat{R}_{{\mathcal{D}}}^{\lambda}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\leq 2R^{\lambda}(f)+2c(B_{0}^{2}+\sigma^{2})\frac{\operatorname{VCdim}({\mathcal{F}})}{n\log^{-1}(n)}$ (26)
where $c$ is an absolute positive constant. Combining (25) and (26) obtains (24). ∎
Proof of Lemma 4.2.
@@
Proof of Theorem 4.3.
According to Lemma 4.2, we set the hypothesis class ${\mathcal{F}}$ as ReQU neural networks ${\mathcal{F}}={\mathcal{N}}(L,S)$ with $L={\mathcal{O}}(\log N)$ and $S={\mathcal{O}}(N^{d})$. Then there exists $f\in{\mathcal{F}}$ such that $\|f-\phi\|_{L^{2}(\mu_{X})}\leq CN^{-s}$ and $\|\nabla f\|_{L^{2}(\nu_{X})}\leq C$. By using Lemma A.6, we find $\operatorname{VCdim}({\mathcal{F}})\leq cSL=cN^{d}\log N$. Substituting these estimates into Lemma 4.1 yields
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}\Big{]}\lesssim CN^{-2s}+C\lambda+(B_{0}^{2}+\sigma^{2})\frac{N^{d}(\log N)^{2}}{n\log^{-1}(n)}.$
Letting $N=(1+\sigma^{2})^{-\frac{1}{d+2s}}n^{\frac{1}{d+2s}}$ and $\lambda=(1+\sigma^{2})^{\frac{2s}{d+2s}}n^{-\frac{2s}{d+2s}}$ deduces the desired result. ∎
Appendix D Proofs of Results in Section 5
Proof of Lemma 5.1.

### v2

@@ Appendix B Proofs of Results in Section III
$|{\mathscr{B}}(f,g)|$ $\leq(\lambda\vee\zeta^{-1/2})\|f\|_{H^{1}(\nu_{X})}\|g\|_{H^{1}(\nu_{X})},$
${\mathscr{B}}(f,f)$ $\geq(\lambda\wedge\kappa^{-1/2})\|f\|_{H^{1}(\nu_{X})},$
for each $f,g\in H^{1}(\nu_{X})$. Further, since that $f_{0}\in L^{2}(\mu_{X})$, the functional $F:H\rightarrow{\mathbb{R}},~{}g\mapsto(f_{0},g)_{L^{2}(\mu_{X})}$ is bounded and linear. Then according to Lax-Milgram theorem [72, Theorem 1 in Chapter 6.2], there exists a unique solution $f^{\lambda}\in H^{1}(\nu_{X})$ to the varitional problem (24). This completes the proof of the uniqueness. See [95, Theorem 2.4.2.7] for the proof of the higher regularity of the solution. ∎
Appendix C Proofs of Results in Section IV
In this section, we demonstrate proofs of theoretical results in Section IV, including Lemma IV.1, Lemma IV.2 and Theorem IV.3. The proof of Lemma IV.1 uses the technique of offset Rademacher complexity, which has been investigated by [96].
Proof of Lemma IV.1.
Recall the population excess risk $R(f)$ and the empirical excess risk $\widehat{R}_{{\mathcal{D}}}(f)$ defined in the proof of Lemma V.2. We further define the regularized excess risk and regularized empirical risk as
$R^{\lambda}(f)$ $:=R(f)+\lambda\|\nabla f\|_{L^{2}(\nu_{X})}^{2},$
$\widehat{R}_{{\mathcal{D}}}^{\lambda}(f)$ $:=\widehat{R}_{{\mathcal{D}}}(f)+\lambda\|\nabla f\|_{L^{2}(\nu_{X})}^{2}.$
@@ Appendix C Proofs of Results in Section IV
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R^{\lambda}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}\lesssim\inf_{f\in{\mathcal{F}}}R^{\lambda}(f)\\ +\frac{B_{0}^{2}+\sigma^{2}}{\log^{-1}n}\inf_{\delta>0}\Big{\{}\frac{2\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))}{n}+\delta\Big{\}}.$ (25)
Before proceeding, we provide the proof sketch. Firstly, in Step (I), we show that
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R^{\lambda}(\widehat{f}_{{\mathcal{D}}}^{\lambda})-2\widehat{R}_{{\mathcal{D}}}^{\lambda}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}$ (26)
$={\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})-2\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}$
$\leq cB_{0}^{2}\inf_{\delta>0}\Big{\{}\frac{2\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))}{n}+\delta\Big{\}},$
where $c$ is an absolute positive constant. It remains to consider the regularized empirical risk. According to (1), we have
$\widehat{L}_{{\mathcal{D}}}^{\lambda}(\widehat{f}_{{\mathcal{D}}}^{\lambda})$
$=\widehat{R}_{{\mathcal{D}}}^{\lambda}(\widehat{f}_{{\mathcal{D}}}^{\lambda})-\frac{2}{n}\sum_{i=1}^{n}\xi_{i}(\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})-f_{0}(X_{i}))+{\mathbb{E}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}\Big{]}.$
Taking expectation with respect to ${\mathcal{D}}\sim\mu^{n}$ on both sides of the equality yields that for each $f\in{\mathcal{F}}$,
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{R}_{{\mathcal{D}}}^{\lambda}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}$
$={\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{L}_{{\mathcal{D}}}^{\lambda}(f)\Big{]}+2{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})\Big{]}-{\mathbb{E}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}\Big{]}$
$\leq R^{\lambda}(f)+\frac{1}{2}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}$
$\quad+c(B_{0}^{2}+\sigma^{2})\inf_{\delta>0}\Big{\{}\frac{2\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))}{n}+\delta\Big{\}},$
@@
$\widehat{R}_{{\mathcal{D}}}^{\lambda}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\leq 2R^{\lambda}(f)\\ +2c(B_{0}^{2}+\sigma^{2})\inf_{\delta>0}\Big{\{}\frac{2\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))}{n}+\delta\Big{\}},$ (27)
where $c$ is an absolute positive constant. Here the inequality invokes
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})\Big{]}\leq\frac{1}{4}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}\\ +\frac{8\sigma^{2}}{n}\inf_{\delta>0}\Big{\{}\frac{\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))}{n}+\delta\Big{\}}\\ +2(B_{0}^{2}+\sigma^{2})\delta,$ (28)
which is obtained in Step (II). Combining (26) and (27) obtains (25).
Step (I). Given a ghost sample ${\mathcal{D}}^{\prime}=\{(X_{i}^{\prime},Y_{i}^{\prime})\}_{i=1}^{n}$, where $\{X_{i}^{\prime}\}_{i=1}^{n}$ are independently drawn from $\mu_{X}$. Further, the ghost sample ${\mathcal{D}}^{\prime}$ is independent of ${\mathcal{D}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$. Let $\varepsilon=\{\varepsilon_{i}\}_{i=1}^{n}$ be a set of Rademacher variables and independent of ${\mathcal{D}}$ and ${\mathcal{D}}^{\prime}$. Since that $\widehat{f}_{{\mathcal{D}}}^{\lambda}\in{\mathcal{F}}$, by the technique of symmetrization, we have
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})-2\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}\leq{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{f\in{\mathcal{F}}}R(f)-2\widehat{R}_{{\mathcal{D}}}(f)\Big{]}$
$={\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{f\in{\mathcal{F}}}{\mathbb{E}}_{{\mathcal{D}}^{\prime}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{2}\Big{]}$
$\quad-\frac{2}{n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{2}\Big{]}$
$\leq{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{{\mathcal{D}}^{\prime}}\Big{[}\sup_{f\in{\mathcal{F}}}\frac{1}{n}\sum_{i=1}^{n}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{2}$
$\quad-\frac{2}{n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{2}\Big{]}$
$={\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{{\mathcal{D}}^{\prime}}\Big{[}\sup_{f\in{\mathcal{F}}}\frac{3}{2n}\sum_{i=1}^{n}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{2}$
$\quad-\frac{1}{2n}\sum_{i=1}^{n}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{2}$
$\quad-\frac{3}{2n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{2}$
$\quad-\frac{1}{2n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{2}\Big{]}$
$\leq{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{{\mathcal{D}}^{\prime}}\Big{[}\sup_{f\in{\mathcal{F}}}\frac{3}{2n}\sum_{i=1}^{n}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{2}$
$\quad-\frac{1}{8B_{0}^{2}n}\sum_{i=1}^{n}(f(X^{\prime}_{i})-f_{0}(X^{\prime}_{i}))^{4}$
$\quad-\frac{3}{2n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{2}$
$\quad-\frac{1}{8B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}\Big{]}$
$={\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\varepsilon}\Big{[}\sup_{f\in{\mathcal{F}}}\frac{3}{n}\sum_{i=1}^{n}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}$
$\quad-\frac{1}{4B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}\Big{]},$ (29)
where the second inequality follows from the convexity of supremum and Jensen’s inequality, and the third inequality is owing to the fact that $0\leq(f(X_{i})-f_{0}(X_{i}))^{2}\leq 4B_{0}^{2}$ for each $f\in{\mathcal{F}}$.
Let $\delta>0$ and let ${\mathcal{F}}_{\delta}$ be an $L^{2}({\mathcal{D}})$ $(B_{0}\delta)$-cover of ${\mathcal{F}}$ satisfying $|{\mathcal{F}}_{\delta}|=N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))$. Then it follows from Cauchy-Schwarz inequality that for each $f\in{\mathcal{F}}$, there exists $f_{\delta}\in{\mathcal{F}}_{\delta}$ such that
$\leq\Big{(}\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})+f_{\delta}(X_{i})-2f_{0}(X_{i}))^{2}$
×$\quad\times(f(X_{i})-f_{\delta}(X_{i}))^{2}\Big{)}^{1/2}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\Big{)}^{1/2}$
$\quad\leq 4B_{0}^{2}\delta.$
By a same argument, we obtain
$\frac{1}{B_{0}^{2}n}\Big{(}-\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}+\sum_{i=1}^{n}(f_{\delta}(X_{i})-f_{0}(X_{i}))^{4}\Big{)}$
$\leq 32B_{0}^{2}\delta.$
Combining (44) with above two inequalities yields
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})-2\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}-20B_{0}^{2}\delta$ (30)
$\leq{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\varepsilon}\Big{[}\max_{f\in{\mathcal{F}}_{\delta}}\frac{3}{n}\sum_{i=1}^{n}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}$
$\quad-\frac{1}{4B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}\Big{]}.$
$\operatorname{Pr}_{\varepsilon}\Big{\{}\frac{3}{n}\sum_{i=1}^{n}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}\\ >t+\frac{1}{4B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}\Big{\}}.$
For a fixed sample ${\mathcal{D}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, the random variables $\{\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}\}_{i=1}^{n}$ are independent and satisfy
${\mathbb{E}}_{\varepsilon}\big{[}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}\big{]}=0,$
and for each $1\leq i\leq n$,
$-(f(X_{i})-f_{0}(X_{i}))^{2}$ $\leq\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}$
$\leq(f(X_{i})-f_{0}(X_{i}))^{2}.$
Consequently, it follows from Hoeffding’s inequality [97, Lemma D.2] that
$\operatorname{Pr}_{\varepsilon}\Big{\{}\frac{3}{n}\sum_{i=1}^{n}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}$
$\quad>t+\frac{1}{4B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}\Big{\}}$
$\leq\exp\Bigg{(}-\frac{(\frac{nt}{3}+\frac{1}{12B_{0}^{2}}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4})^{2}}{2\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}}\Bigg{)}$
$\leq\exp\Big{(}-\frac{nt}{18B_{0}^{2}}\Big{)},$
${\mathbb{E}}_{\varepsilon}\Big{[}\max_{f\in{\mathcal{F}}_{\delta}}\frac{3}{n}\sum_{i=1}^{n}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}$
$\quad-\frac{1}{4B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}\Big{]}$
[…]
$=T+\frac{18B_{0}^{2}}{n}N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))\exp\Big{(}-\frac{nT}{18B_{0}^{2}}\Big{)}.$
By setting $T=18B_{0}^{2}\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))n^{-1}$, we deduces
${\mathbb{E}}_{\varepsilon}\Big{[}\max_{f\in{\mathcal{F}}_{\delta}}\frac{3}{n}\sum_{i=1}^{n}\varepsilon_{i}(f(X_{i})-f_{0}(X_{i}))^{2}$ (31)
$\quad-\frac{1}{4B_{0}^{2}n}\sum_{i=1}^{n}(f(X_{i})-f_{0}(X_{i}))^{4}\Big{]}$
$\leq\frac{18B_{0}^{2}}{n}(1+\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))).$
Combining (30) and (31) implies that
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}R(\widehat{f}_{{\mathcal{D}}}^{\lambda})-2\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}$ (32)
$\leq\frac{18B_{0}^{2}}{n}(1+\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}})))+20B_{0}^{2}\delta.$
This completes the proof of (26).
Step (II). Recall the $L^{2}({\mathcal{D}})$ $(B_{0}\delta)$-cover ${\mathcal{F}}_{\delta}$ of the hypothesis class ${\mathcal{F}}$. There exists $f_{\delta}\in{\mathcal{F}}_{\delta}$ such that
$\frac{1}{n}\sum_{i=1}^{n}|f_{\delta}(X_{i})-\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})|^{2}\leq(B_{0}\delta)^{2},$
which implies
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}(\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})-f_{\delta}(X_{i}))\Big{]}$ (33)
$\leq{\mathbb{E}}_{{\mathcal{D}}}^{1/2}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}\Big{]}{\mathbb{E}}_{{\mathcal{D}}}^{1/2}\Big{[}\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})-f_{\delta}(X_{i}))^{2}\Big{]}$
$\leq B_{0}\sigma\delta,$
and
$\widehat{R}_{{\mathcal{D}}}^{1/2}(f_{\delta})$ (34)
$\leq\Big{(}\frac{1}{n}\sum_{i=1}^{n}(f_{\delta}(X_{i})-\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i}))^{2}\Big{)}^{1/2}+\widehat{R}_{{\mathcal{D}}}^{1/2}(\widehat{f}_{{\mathcal{D}}}^{\lambda})$
$\leq B_{0}\delta+\widehat{R}_{{\mathcal{D}}}^{1/2}(\widehat{f}_{{\mathcal{D}}}^{\lambda}),$
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})\Big{]}$
$={\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}(\widehat{f}_{{\mathcal{D}}}^{\lambda}(X_{i})-f_{0}(X_{i}))\Big{]}$
$\leq{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\xi_{i}(f_{\delta}(X_{i})-f_{0}(X_{i}))\Big{]}+B_{0}\sigma\delta$
@@
$\leq{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\frac{\widehat{R}_{{\mathcal{D}}}^{1/2}(\widehat{f}_{{\mathcal{D}}}^{\lambda})+B_{0}\delta}{\sqrt{n}}\psi(f_{\delta})\Big{]}+B_{0}\sigma\delta$
$\leq\Big{(}{\mathbb{E}}_{{\mathcal{D}}}^{1/2}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+B_{0}\delta\Big{)}\frac{1}{\sqrt{n}}{\mathbb{E}}_{{\mathcal{D}}}^{1/2}\Big{[}\psi^{2}(f_{\delta})\Big{]}+B_{0}\sigma\delta$
$\leq\frac{1}{4}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}}}^{\lambda})\Big{]}+\frac{2}{n}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\psi^{2}(f_{\delta})\Big{]}$
$\quad+\frac{1}{4}B_{0}^{2}\delta^{2}+B_{0}\sigma\delta.$ (35)
${\mathbb{E}}_{\xi}\Big{[}\psi^{2}(f_{\delta})\Big{]}\leq{\mathbb{E}}_{\xi}\Big{[}\max_{f\in{\mathcal{F}}_{\delta}}\psi^{2}(f)\Big{]}\leq 4\sigma^{2}(\log|{\mathcal{F}}_{\delta}|+1).$ (36)
Combining (35) and (36) yields (28). ∎
Proof of Lemma IV.2.
@@
Proof of Theorem IV.3.
According to Lemma 4.2, we set the hypothesis class ${\mathcal{F}}$ as ReQU neural networks ${\mathcal{F}}={\mathcal{N}}(L,S)$ with $L={\mathcal{O}}(\log N)$ and $S={\mathcal{O}}(N^{d})$. Then there exists $f\in{\mathcal{F}}$ such that $\|f-\phi\|_{L^{2}(\mu_{X})}\leq CN^{-s}$ and $\|\nabla f\|_{L^{2}(\nu_{X})}\leq C$. By using Lemma A.5 and set $\delta=1/n$, we find
$\log N(B_{0}n^{-1},{\mathcal{F}},L^{2}({\mathcal{D}}))\\ \lesssim LS\log S\log n\lesssim N^{d}\log^{2}N\log n.$
Substituting these estimates into Lemma IV.1 yields
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}\Big{]}\\ \lesssim CN^{-2s}+C\lambda+C\log n\frac{N^{d}\log^{2}N}{n\log^{-1}(n)}.$
Letting $N={\mathcal{O}}(n^{\frac{1}{d+2s}})$ and $\lambda={\mathcal{O}}(n^{-\frac{2s}{d+2s}}\log^{3}n)$ deduces the desired result. ∎
Appendix D Proofs of Results in Section V
In this section, we show proofs of theoretical results in Section V. The proofs for the deep Sobolev regressor are shown in Section D-A, and proofs for semi-supervised deep Sobolev regressor are shown in Section D-B.
Proof of Lemma V.1.